OpenStudy (anonymous):

What is the solution of the equation 9 = -2x + 15? A. -21 B. -12 C. 6 D. 12

OpenStudy (luigi0210):

First move the 15 over: 9-15=-2x -6=-2x Now divide: 3=x

OpenStudy (anonymous):

A,B,C,or D

OpenStudy (texaschic101):

9 = -2x + 15 9 - 15 = - 2x -6 = - 2x 3 = x check... 9 = - 2x + 15 9 = -2(3) + 15 9 = -6 + 15 9 = 9 (correct) Are you sure you wrote the equation correctly, or are your answer choices written correctly. Because the answer is 3

OpenStudy (e.cociuba):

^^ he's right the answer is 3. Hmm maybe the multiply choice is wrong or maybe u made a mistake.

OpenStudy (anonymous):

i did not make a mistake and the multiply choice is not wrong

OpenStudy (e.cociuba):

Even this site says the answers 3 :) http://www.wolframalpha.com/input/?i=What+is+the+solution+of+the+equation++9+%3D+-2x+%2B+15%3F

OpenStudy (anonymous):

it is B

OpenStudy (e.cociuba):

how so?

OpenStudy (texaschic101):

you think it is B.....lets check.. 9 = - 2x + 15 9 = -2(-12) + 15 9 = 24 + 15 9 = 39 (I don't think so)